A május 2-i dolgozat eredményei.

feladatok  1.  2.  3.  4.  5  6.

diákok:

Andrékó Judit + +
Balogh Péter  +
Benczik Attila  1/2  +
Czibere Marianna  +  1/2 1/2
Dienes Zoltán  1/2  +  +
Fórizs Fruzsina  +  1/2  1/2  +
Gombkötő Angéla  1/2
Iván Éva  1/2
Kerékgyártó Zita  +  +  3/4 +  +
– Kocsis Zsuzsanna  +  +
Orosz Erzsébet  +  +  +  +
Pintér Györgyi  +  +  1/2
Szilágyi Renáta Judit  +  +
Sztankó Paticia  1/2  1/2  1/2 +  +